13.3 Estimate the amount of solar heating that is lost as a result of using one or two tempered 1/8 in thick glass plates covBoth please
Answer #1

a) The solar heat is produced by Infrared radiation which covers the spectrum form 700nm to 1mm. For wavelength range 200nm to 3000nm, the transmission coefficient is 80% and for wavelengths greater than 3000nm, transmission coefficient is 2%.

The 700nm to 3000nm band forms %age of total IR band:

\frac{(3000-700)nm}{1mm-700nm} \approx \frac{2300 \times 10^{-9}m}{10^{-3}m}=0.23 \times 10^{-2}

Thus, 0.23% of IR band has transmission coefficient 80%.

The rest 99.77% has transmission coefficient 2%. So, the total transmission:

0.23 \times 0.8 +99.77 \times 0.02=0.184+1.9954=2.18

So, the total transmission is 2.18%. This much amount of heat escapes through glass plates.

b) The change in Q_s if any of its variable changes :

dQ_s=\frac{\partial Q_s}{\partial A_c}dA_c+\frac{\partial Q_s}{\partial S}dS

Q_s=L\left [ 0.8+ln\left \{ \left ( A_C/L \right )^{2/3} \left ( S/L \right )^{1/2} \right \} \right ]

It can be further simplified using properties of logarithmic functions:

Q_s=L\left [ 0.8+\frac{2}{3}ln\left ( A_c/L \right )+\frac{1}{2}ln(S/L) \right ]

Calculating the partial derivatives:

\frac{\partial Q_s}{\partial A_c}=\frac{2}{3} \frac{1}{A_c/L} \times 1/L=\frac{2}{3A_c}

\frac{\partial Q_s}{\partial S}=\frac{1}{2} \frac{1}{S/L} \times 1/L=\frac{1}{2S}

Substituting the values, we get

dQ_s=\frac{2}{3A_c}dA_c+\frac{1}{2 S}dS

If collector area is increased by 10%, then

dA_c=0.1

Then the values of Q_s changes by:

\frac{2}{3} \times 0.1=0.07

That it, 7%.

If storage is increased by 10%, then change in Q_s:

\frac{1}{2} \times 0.1=0.05 i.e. 5%